Write a proportion. Then determine the missing measure.
1. CHIMNEYS A 6-ft observer casts a 4-ft. shadow at the same time the chimney on the Ohio Power Company casts an 804-foot shadow. How tall is the chimney?
2. BUILDINGS The May Road Apartments in Hong Kong cast a 90-meter shadow at the same time a 1.5-meter tall tenant casts a 0.75-meter shadow. How tall is the apartment building?
3. WORLD RECORDS The world's tallest man lived from 1918 to 1940. he cast a
4-foot 5 1/2inch shadow when a 6-foot pole cast a 3-foot shadow.
How tall was he?
4. SHADOWS A man casts a 14-foot shadow. A 4-foot child casts a 9-foot 4-inch shadow at the same time. How tall is the man?

Answer by checkley75(3666)

4/6=408/X CROSS MULTIPLY
4X=408*6
4X=2448
X=2448/4
X=612 FEET TALL FOR THE CHIMNEY
-------------------------------------------------------------
90/X=.75/1.5
.75X=90*1.5
.75X=135
X=135/.75
X=180 METERS TALL
------------------------------------------------------------
4FT 5 1/2 IN/X=3/6
3X=6*53.5
3X=321
X=321/3
X=107 INCHES
------------------------------------------------------------
14/X=9FT4IN/4
112X=14*12*4
112X=672
X=672/112
X=6 FT. TALL FOR THE MAN.